OpenStudy (anonymous):

Which expression represents the quotient? 8x^6z^4+4x^4z^2/4x^2z A.2x^4z^3+x^2z B.2x^3z^4+x^2z^2 C.4x^4z^3+3x^2z D.4x^3z^4+3x^2z^2

OpenStudy (mimi_x3):

@bibby I leave the dirty work for you. Pls go ahead

OpenStudy (bibby):

\(\huge\frac{a+b+c}{d}=\frac{a}{d}+\frac{b}{d}+\frac{c}{d}\) \(\huge\frac{a^m}{a^n}=a^{(m-n)}\)

OpenStudy (mimi_x3):

Thanks bby <3

OpenStudy (bibby):

mhm ;*